OpenStudy (anonymous):

x^4-7x^3+14x^2-38x-60, zero given:1+3i I gotten that factored (x-(1-3i)(x-(1+3i) then got (x^2-2x+10). I don't know what to factor with it

OpenStudy (perl):

there is an important rule complex zeroes come in conjugate pairs, always. so if 1 +3i is a zero, then so is (1 - 3i )

OpenStudy (anonymous):

Yeah I understand that

OpenStudy (perl):

so first expand ( x - (1 + 3i)) ( x - ( 1 - 3i))

OpenStudy (perl):

( x - (1 + 3i)) ( x - ( 1 - 3i)) = ( x - 1 - 3i ) ( x - 1 + 3i ) = ( (x-1) - 3i) ( (x-1) + 3i ) = ( x-1)^2 - 3i*(x-1) + 3i(x-1) - 3i^2

OpenStudy (perl):

with me so far

OpenStudy (perl):

= (x-1)^2 -3*-1 = x^2 -2x + 1 + 9 = x^2 -2x + 10

OpenStudy (perl):

now do long division

OpenStudy (perl):



OpenStudy (anonymous):

I don't understand what you did in the third line. I multiplied the second line together and that's how I got x^2-2x+10

OpenStudy (perl):

that was just a trick to rearrange parentheses, it doesn't matter. we got the same answer

OpenStudy (anonymous):

Oh okay, then I don't know what I would factor with that to get back to the original polynomial with that

OpenStudy (perl):